OpenStudy (ashy98):

Can someone help me??(:

OpenStudy (ashy98):

\[What is product of 5\sqrt{3}\times 4 \sqrt{15}? simplify if possible.\] \[A.) 60\sqrt{15}\] \[B.) 20\sqrt{5}\] \[C.) 20\sqrt{15}\] \[D.) 60\sqrt{5}\]

OpenStudy (ashy98):

@Teddyiswatshecallsme @amistre64 @KyanTheDoodle

OpenStudy (amistre64):

what have you attemoted?

OpenStudy (amistre64):

**attempted

OpenStudy (ashy98):

c? but im really not sure

OpenStudy (amistre64):

C is like totally not it

OpenStudy (amistre64):

sqrt(a) * sqrt(b) = sqrt(ab)

OpenStudy (amistre64):

so even tho 5*4 = 20, 3*15 isnt 15

OpenStudy (ashy98):

so it would be B?

OpenStudy (amistre64):

if 3*15 isnt 15, then it sure isnt goinna be 5 it aint B

OpenStudy (amistre64):

5*4 = 20 sqrt(3)*sqrt(15) = sqrt(???)

OpenStudy (ashy98):

45

OpenStudy (amistre64):

good so the solution is some simplified version of 20 sqrt(45) can sqrt(45) be simplified? any perfect square factors in 45?

OpenStudy (amistre64):

maybe 9*5 = 45?

OpenStudy (amistre64):

20 sqrt(45) 20 sqrt(9*5) 20 sqrt(9) sqrt(5)

OpenStudy (ashy98):

so that would be D?

OpenStudy (amistre64):

60 sqrt(5) yes :)

OpenStudy (ashy98):

yay thank you for helping me!(: